1. a. In the laboratory, a general chemistry student measured the pH of a 0.331 M aqueous solution of hypochlorous acid to be 3.986.
Use the information she obtained to determine the Ka for this acid.

b.In the laboratory, a general chemistry student measured the pH of a 0.331 M aqueous solution of acetylsalicylic acid (aspirin), HC9H7O4 to be 1.985.
Use the information she obtained to determine the Ka for this acid.
Answer #1

Holl dress ociutes in water as- И01 (4) — что он ( ) ( ) я. 3 - 986 Lя*] - В-936) ... [+) = ). 032 x 4 c ) [] - [ ] - - 32 yoТІ а Suppete 4 не и i к». ktopң Aspirin dissacrates us- Воон — M Reoo + Тач) (0) ри. 1983 Ск? - (- 1-985) [1+1 = 0 - 0 \о 35